OpenStudy (anonymous):

Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have? I got this so far: d+q=46 I don't know what the second problem should be.

OpenStudy (anonymous):

yeah your equation is correct d+q = 46 which gives d = 46-q now use the dollar value given d*10 + q*25 = 700 so (46-q)*10 + q*25 = 700 Solve for q get d from q